# Volume-law entanglement fragmentation of quasiparticles

### v1

@@ II.3 States with mixed occupancy
$z_{i}\equiv\lim_{L\to+\infty}\frac{p_{i}}{L}.$ (24)
@@
For example, we may have the unit pattern with $l_{1}=2$ and $\kappa_{1}=\{0\}$ excited in the momentum space $[0,\frac{L}{3}-1]$, and the unit pattern with $l_{2}=3$ and $\kappa_{2}=\{0,1\}$ excited in the momentum space $[\frac{L}{2},\frac{3L}{4}-1]$. For $L=24$, such a state has the excited modes
$K=(3\kappa_{1})\bigcup(2\kappa_{2})=\{0,2,4,6,12,13,15,16\},$ (25)
@@
$S_{L,xL,\bigcup_{i=1}^{r}(p_{i}\kappa_{i})}^{(n),\text{fer}}\approx\sum_{i=1}^{r}p_{i}s_{l_{i},\kappa_{i}}^{(n),\text{fer}}(x),$ (27)
@@
$\lim_{L\to+\infty}\frac{S_{L,xL,\bigcup_{i=1}^{r}(p_{i}\kappa_{i})}^{(n),\text{fer}}}{L}=\sum_{i=1}^{r}z_{i}s_{l_{i},\kappa_{i}}^{(n),\text{fer}}(x).$ (28)
This is the main result of this paper. We have expressed the volume-law many-body entanglement entropy as the sum of few-body entanglement entropies. It is easy to confirm the formula (28) numerically. The results are also the same as those in the limit $L_{A}\ll L$ found in Alba:2009th .
III Interacting fermions
The numerical verification of formulas (22) and (28) in free fermionic chains motivates us to further conjecture that these formulas still hold true in integrable interacting fermionic chains. By applying the Jordan-Wigner transformation, the transverse field Ising chain, given by
$H=-\frac{1}{2}\sum_{j=1}^{L}\big{(}\sigma_{j}^{x}\sigma_{j+1}^{x}+h\sigma_{j}^{z}\big{)},$ (29)
is transformed into a fermionic chain with nearest-neighbor interactions. This model is solvable, as shown in Lieb:1961fr ; Pfeuty:1970ayt , and the entanglement entropy can be calculated following the methods outlined in Vidal:2002rm ; Latorre:2003kg ; Alba:2009th . We have carried out extensive numerical verifications of formulas (22) and (28) in the Ising chain, thereby confirming the aforementioned conjecture.
IV Spin-1/2 XXZ chain
@@
$H=-\frac{1}{4}\sum_{j=1}^{L}\big{(}\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+\Delta\sigma_{j}^{z}\sigma_{j+1}^{z}\big{)},$ (30)
with periodic boundary conditions for the Pauli matrices $\sigma_{L+1}^{\alpha}=\sigma_{1}^{\alpha}$, $\alpha=x,y,x$. The excited states in the XXZ chain can be obtained from the coordinate Bethe ansatz Karbach:1998abi ; Karbach:1998oja ; Karbach:2000ngx ; Caux:2014uuq .
We consider the energy eigenstates $|\mathcal{I}\rangle$ with an extensive number of magnons, which are labeled by the Bethe numbers =$\mathcal{I}=\{I_{1},I_{2},\cdots\}$. For example, there may be a length $l$ unit pattern $\iota=\{0\}$ excited in the whole or part of the space of the Bethe numbers $[0,L-1]$, with the number of repetitions being $p$. When $\Delta=2$, this corresponds to a length $l-1$ unit pattern $\kappa=\{0\}$ with repetition $p$ in the momentum space. We obtain the Rényi entropy and entanglement entropy
$S^{(n),\text{XXZ}}_{L,xL,p\iota}\approx ps^{(n),\text{fer}}_{l-1,\kappa}(x),$ (31)
@@
Figure 3: Entanglement entropy in states with an extensive number of magnons in the XXZ chain (symbols) is compared with the corresponding analytical formula from the fermionic chain (blue dashed lines).
V Free bosons
@@
$H_{\text{bos}}=\sum_{j=1}^{L}\left(a_{j}^{\dagger}a_{j}+\frac{1}{2}\right),$ (32)
@@
with the bosonic modes $a_{j}$ and $a_{j}^{\dagger}$ satisfying $[a_{j_{1}},a_{j_{2}}]=[a^{\dagger}_{j_{1}},a^{\dagger}_{j_{2}}]=0$ and $[a_{j_{1}},a^{\dagger}_{j_{2}}]=\delta_{j_{1}j_{2}}$. In an energy eigenstate, the entanglement entropy can be obtained from the subsystem mode method Zhang:2021bmy , and the Rényi entropy could be obtained from the permanent formula derived in Zhang:2020dtd using the wavefunction method Bombelli:1986rw ; Srednicki:1993im ; Castro-Alvaredo:2018dja ; Castro-Alvaredo:2018bij .
@@
We consider the entanglement entropy in a state with repeated unit patterns in the bosonic chain. Following (13) in the fermionic chain, we define
$s^{(n),\text{bos}}_{l,\kappa}(x)\equiv\lim_{p\to+\infty}\frac{S_{pl,xpl,p\kappa}^{(n),\text{bos}}}{pl}.$ (33)
@@
Following (14), one might naively expect that the function $s^{(n),\text{bos}}_{l,\kappa}(x)$ takes the form
$s_{l,\kappa}^{(n),\text{naive}}(x)\equiv yS_{l,\alpha+1,\kappa}^{(n),\text{bos}}+(1-y)S_{l,\alpha,\kappa}^{(n),\text{bos}},$ (34)
@@
with the definitions of $\alpha$ and $y$ given in (II.1). However, numerical results show that while the function $s^{(n),\text{bos}}_{l,\kappa}(x)$ is well-defined, it differs from the naive expectation $s_{l,\kappa}^{(n),\text{naive}}(x)$. We present numerical examples in Fig. 4.
@@
Figure 4: The entanglement entropy in free bosonic chains (symbols) is contrasted with the naive expectation $s_{l,\kappa}^{(n),\text{naive}}(x)$ (34) (red dashed lines). In each panel, the blue dashed line represents the interpolation of the results corresponding to the maximum $L$.
VI Interacting bosons
@@
Similarly to (28) in the fermionic chain, we conjecture that in the free and interacting bosonic chains, there is
$\lim_{L\to+\infty}\frac{S_{L,xL,\bigcup_{i=1}^{r}(p_{i}\kappa_{i})}^{(n),\text{bos}}}{L}=\sum_{i=1}^{r}z_{i}s_{l_{i},\kappa_{i}}^{(n),\text{bos}}(x),$ (35)
with the definition of $z_{i}$ in (24) and the definition of $s_{l_{i},\kappa_{i}}^{(n),\text{bos}}(x)$ in (33). We consider the harmonic chain
$H=\frac{1}{2}\sum_{j=1}^{L}\left[p_{j}^{2}+m^{2}q_{j}^{2}+(q_{j}-q_{j+1})^{2}\right],$ (36)
@@
with the periodic boundary condition $q_{L+1}=q_{j}$ and the commutation relations $[q_{j_{1}},q_{j_{2}}]=[p_{j_{1}},p_{j_{2}}]=0$ and $[q_{j_{1}},p_{j_{2}}]=i\delta_{j_{1}j_{2}}$. The harmonic chain is essentially a chain of bosons with nearest-neighbor interactions. The above conjecture could be easily checked by numerically calculating the Rényi entropy in excited energy eigenstates following the wavefunction method Castro-Alvaredo:2018dja ; Castro-Alvaredo:2018bij ; Zhang:2020ouz ; Zhang:2020txb .
VII Discussions
@@
We have studied the entanglement entropy in quasiparticle states that exhibit repeated and sequential unit patterns in momentum space. We found that each unit pattern contributes independently and universally to the entanglement in the scaling limit, leading to a volume-law scaling of entanglement entropy. Numerical simulations confirmed these findings, demonstrating volume-law entanglement fragmentation in both fermionic and bosonic chains. We derived an analytical formula for free fermions that is applicable not only to interacting fermions in the scaling limit but also, with certain identifications, to the spin-1/2 XXZ chain.
@@
It was suggested in Castro-Alvaredo:2018dja to harness the entanglement of quasiparticles for quantum information purposes. However, the entanglement among a small number of quasiparticles is delicate and susceptible to environmental disruptions. The results presented in this paper imply that by successively exciting repeated patterns in momentum space, it is possible to create a robust volume-law entanglement of quasiparticles.
@@ Acknowledgements
The author thanks Olalla A. Castro-Alvaredo and M. A. Rajabpour for reading a previous versions of the draft and helpful comments and discussions. JZ acknowledgements support from the National Natural Science Foundation of China (NSFC) grant number 12205217.

### v2

@@ 2.3 States with mixed occupancy
$z_{i}\equiv\lim_{L\to+\infty}\frac{p_{i}}{L}.$ (2.25)
@@
For example, we may have the unit pattern with $l_{1}=2$ and $\kappa_{1}=\{0\}$ excited in the momentum space $0,\frac{L}{3}-1]$, and the unit pattern with $l_{2}=3$ and $\kappa_{2}=\{0,1\}$ excited in the momentum space $\frac{L}{2},\frac{3L}{4}-1]$. For $L=24$, such a state has the excited modes
$K=(3\kappa_{1})\bigcup(2\kappa_{2})=\{0,2,4,6,12,13,15,16\},$ (2.26)
@@
$S_{L,xL,\bigcup_{i=1}^{r}(p_{i}\kappa_{i})}^{(n),\text{fer}}\approx\sum_{i=1}^{r}p_{i}s_{l_{i},\kappa_{i}}^{(n),\text{fer}}(x),$ (2.28)
@@
$\lim_{L\to+\infty}\frac{S_{L,xL,\bigcup_{i=1}^{r}(p_{i}\kappa_{i})}^{(n),\text{fer}}}{L}=\sum_{i=1}^{r}z_{i}s_{l_{i},\kappa_{i}}^{(n),\text{fer}}(x).$ (2.29)
This is the main result of this paper. We have expressed the volume-law many-body entanglement entropy as the sum of few-body entanglement entropies. It is easy to confirm the formula (2.29) numerically. The results are also the same as those in the limit $L_{A}\ll L$ found in [15]. We show examples of the results in Fig. 4.
Figure 4: The many-body entanglement entropy in the states with partial occupation (symbols) and the result from the few-body entanglement entropy (blue dashed lines). The unit patterns $l_{1},\kappa_{1}$, $l_{2},\kappa_{2}$ and their corresponding repetition ratios $z_{1}$, $z_{2}$ are shown in each panel. On the left panel the two different patterns are adjacent in the momentum space, while on the right panel the two patterns are disjoint.
The results indicate the effective form of the RDM
$\rho_{L,L_{A},\bigcup_{i=1}^{r}(p_{i}\kappa_{i})}\sim\Big{[}\bigotimes_{i=1}^{r}\Big{(}\underbrace{\rho_{l_{i},\alpha_{i}+1,\kappa_{i}}\otimes\cdots\otimes\rho_{l_{i},\alpha_{i}+1,\kappa_{i}}}_{\lfloor z_{i}a_{i}\rfloor}\otimes\underbrace{\rho_{l_{i},\alpha_{i},\kappa_{i}}\otimes\cdots\otimes\rho_{l_{i},\alpha_{i},\kappa_{i}}}_{\lfloor z_{i}(\lfloor\frac{L}{l_{i}}\rfloor-a_{i})\rfloor}\Big{)}\Big{]}\otimes\rho^{\prime}_{\lceil(1-z)L_{A}\rceil},$ (2.30)
where $\alpha_{i}\equiv\lfloor xl_{i}\rfloor$$a_{i}\equiv(L_{A}{\rm~mod~}\lfloor\frac{L}{l_{i}}\rfloor)$, and $\rho^{\prime}_{\lceil(1-z)L_{A}\rceil}$ is a density matrix of about $\lceil(1-z)L_{A}\rceil$ sites with sub-volume-law entanglement entropy. We have also defined $z=\sum_{i=1}^{r}z_{i}$.
3 Interacting fermions
The numerical verification of formulas (2.22) and (2.29) in free fermionic chains motivates us to further conjecture that these formulas still hold true in integrable interacting fermionic chains. We consider the the transverse field Ising chain
$H=-\frac{1}{2}\sum_{j=1}^{L}\big{(}\sigma_{j}^{x}\sigma_{j+1}^{x}+h\sigma_{j}^{z}\big{)}.$ (3.1)
We use $j=1,2,\cdots,L$ to label the sites of the chain. On each site, there is a qubit represented by Pauli matrices $\sigma_{j}^{x},\sigma_{j}^{y},\sigma_{j}^{z}$, and we impose periodic boundary conditions such that $\sigma_{L+1}^{x}=\sigma_{1}^{x}$. By applying the Jordan-Wigner transformation, the chain is transformed into a fermionic chain with nearest-neighbor interactions. This model is solvable, as shown in [55, 56], and the entanglement entropy can be calculated following the methods outlined in [10, 11, 15]. Although we cannot derive the analytical form of the entanglement entropy in such a model, we have carried out extensive numerical verifications of formulas (2.22) and (2.29) in the Ising chain, thereby confirming the aforementioned conjecture. We show examples of numerical results in Fig. 5.
Figure 5: The many-body entanglement entropy in interacting fermions (symbols) and the corresponding analytical results from the few-body entanglement entropy (blue dashed lines). On the left panel, there is one pattern that occupies the whole momentum space. On the right panel, there are two adjacent patterns, with one occupying 1/3 of the momentum space and the other occupying 1/2 of the momentum space. We have used $h=1$ for both panels.
4 Spin-1/2 XXZ chain
@@
$H=-\frac{1}{4}\sum_{j=1}^{L}\big{(}\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+\Delta\sigma_{j}^{z}\sigma_{j+1}^{z}\big{)},$ (4.1)
with periodic boundary conditions for the Pauli matrices $\sigma_{L+1}^{\alpha}=\sigma_{1}^{\alpha}$, $\alpha=x,y,x$. The excited states in the XXZ chain can be obtained from the coordinate Bethe ansatz [57, 58, 59, 60].
We consider the energy eigenstates $|\mathcal{I}\rangle$ with an extensive number of magnons, which are labeled by the Bethe numbers $\mathcal{I}=\{I_{1},I_{2},\cdots\}$. For example, there is a length $l$ unit pattern $\iota=\{0\}$ excited in the whole or part of the space of the Bethe numbers [0,L1][0,L-1], with the number of repetitions being $p$. When $\Delta=2$, this corresponds to a length l1l-1 unit pattern $\kappa=\{0\}$ with repetition $p$ in the momentum space. We obtain the Rényi entropy and entanglement entropy
$S^{(n),\text{XXZ}}_{L,xL,p\iota}\approx ps^{(n),\text{fer}}_{l-1,\kappa}(x),$ (4.2)
@@
Figure 6: Entanglement entropy in states with an extensive number of magnons in the XXZ chain (symbols) is compared with the corresponding analytical formula from the fermionic chain (blue dashed lines). On the left panel, there is one pattern that occupies the whole momentum space. On the right panel, there is one pattern that occupies 1/2 of the momentum space. We have used $\Delta=2$ for both panels.
5 Free bosons
@@
$H_{\text{bos}}=\sum_{j=1}^{L}\left(a_{j}^{\dagger}a_{j}+\frac{1}{2}\right),$ (5.1)
@@
with the bosonic modes $a_{j}$ and $a_{j}^{\dagger}$ satisfying $a_{j_{1}},a_{j_{2}}]=[a^{\dagger}_{j_{1}},a^{\dagger}_{j_{2}}]=0$ and $a_{j_{1}},a^{\dagger}_{j_{2}}]=\delta_{j_{1}j_{2}}$. In an energy eigenstate, the entanglement entropy can be obtained from the subsystem mode method [33], and the Rényi entropy with fixed index $n$ could be obtained from the permanent formula derived in [32] using the wavefunction methods [61, 62, 25, 26].
@@
We consider the entanglement entropy in a state with repeated unit patterns in the bosonic chain. Following (2.13) in the fermionic chain, we define the finite contributions to the entanglement from each unit pattern
$s^{(n),\text{bos}}_{l,\kappa}(x)\equiv\lim_{p\to+\infty}\frac{S_{pl,xpl,p\kappa}^{(n),\text{bos}}}{p}.$ (5.2)
@@
Following (2.14), one might naively expect that the function $s^{(n),\text{bos}}_{l,\kappa}(x)$ takes the form
$s_{l,\kappa}^{(n),\text{naive}}(x)\equiv yS_{l,\alpha+1,\kappa}^{(n),\text{bos}}+(1-y)S_{l,\alpha,\kappa}^{(n),\text{bos}},$ (5.3)
@@
with the definitions of $\alpha$ and $y$ given in (2.1). However, numerical results show that while the function $s^{(n),\text{bos}}_{l,\kappa}(x)$ is well-defined, it differs from the naive expectation $s_{l,\kappa}^{(n),\text{naive}}(x)$. We present numerical examples in Fig. 7. It remains an open problem to derive the explicit analytical expression of $s^{(n),\text{bos}}_{l,\kappa}(x)$. The bosonic version of the formulas (2.22) and (2.29) still apply with the definitions (5.2), and we have verified the formulas with numerically examples, which are not shown here.
@@
Figure 7: The entanglement entropy in free bosonic chains (symbols) is contrasted with the naive expectation $s_{l,\kappa}^{(n),\text{naive}}(x)$ (5.3) (red dashed lines). In each panel, the blue dashed line represents the interpolation of the results corresponding to the maximum $L$, i.e. $L=24$ for panel (a) and $L=18$ for panel (b). In each panel, there is one pattern that occupies the whole momentum space.
6 Interacting bosons
@@
Similarly to (2.29) in the fermionic chain, we conjecture that in the free and interacting bosonic chains, there is
$\lim_{L\to+\infty}\frac{S_{L,xL,\bigcup_{i=1}^{r}(p_{i}\kappa_{i})}^{(n),\text{bos}}}{L}=\sum_{i=1}^{r}z_{i}s_{l_{i},\kappa_{i}}^{(n),\text{bos}}(x),$ (6.1)
with the definition of $z_{i}$ in (2.25) and the definition of $s_{l_{i},\kappa_{i}}^{(n),\text{bos}}(x)$ in (5.2). We consider the harmonic chain
$H=\frac{1}{2}\sum_{j=1}^{L}\left[p_{j}^{2}+m^{2}q_{j}^{2}+(q_{j}-q_{j+1})^{2}\right],$ (6.2)
@@
with the periodic boundary condition $q_{L+1}=q_{j}$ and the commutation relations $q_{j_{1}},q_{j_{2}}]=[p_{j_{1}},p_{j_{2}}]=0$ and $q_{j_{1}},p_{j_{2}}]=i\delta_{j_{1}j_{2}}$. The harmonic chain is essentially a chain of bosons with nearest-neighbor interactions. The above conjecture could be easily checked by numerically calculating the Rényi entropy in excited energy eigenstates following the wavefunction methods [25, 26, 63, 64]. In Fig. 8, we provide numerical results supporting the conjecture. For larger masses $m$, the model approaches free bosons and exhibits smaller numerical errors. Conversely, smaller masses $m$ lead to larger numerical errors.
Figure 8: The many-body entanglement entropy in interacting bosons (symbols) and the results $s_{l_{1},\kappa_{1}}^{(2)}(x)$ and $zs_{l_{1},\kappa_{1}}^{(2)}(x)$ (blue lines) with $l_{1}=2,\kappa_{1}=\{0\}$ and $s_{l,\kappa}^{(2)}(x)$ taking the approximate value from the interpolation in panel (a) of Fig. 7. On the top panels, there is one pattern that occupies the whole momentum space. On the bottom panels, there is one pattern that occupies 1/2 of the momentum space. In different columns, we have used different values of $m$.
7 Discussions
@@
We have studied the entanglement entropy in quasiparticle states that exhibit repeated and sequential unit patterns in momentum space. We found that each unit pattern contributes independently and universally to the entanglement in the scaling limit, leading to a volume-law scaling of entanglement entropy. Numerical simulations confirmed these findings, demonstrating volume-law entanglement fragmentation in both fermionic and bosonic chains. We derived an analytical formula for free fermions that is applicable not only to interacting fermions in the scaling limit but also, with certain identifications, to the spin-1/2 XXZ chain. Numerical results of free and interacting bosonic chains reveal universal entanglement characteristics in bosonic systems; however, explicit analytical formulations remain elusive. Deriving such analytical expressions for bosonic chains would be an important direction for future work.
The results in this paper reveal a fundamental distinction between fragmented volume-law entanglement and thermal volume-law entanglement. Thermal volume-law entanglement, found in states such as random states or eigenstates of chaotic Hamiltonians [37, 38, 47], emerges from maximal scrambling and adherence to the eigenstate thermalization hypothesis (ETH), resulting in featureless, homogeneous entanglement devoid of substructure. In contrast, fragmented volume-law entanglement in quasiparticle states arises from repeated unit patterns in momentum space and decomposes the state into a sum of few-body entangled components. Although both types exhibit extensive entanglement, fragmented states violate ETH due to the underlying integrability, contrasting sharply with the thermal case. The existence and universality of entanglement fragmentation remain to be further studied. We conjecture that this phenomenon may stem from the existence and independence of quasiparticles resulting from integrability.
@@
It was suggested in [25] to harness the entanglement of quasiparticles for quantum information purposes. However, the entanglement among a small number of quasiparticles is delicate and susceptible to environmental disruptions. The results presented in this paper imply that by successively exciting repeated patterns in momentum space, it is possible to create a robust volume-law entanglement of quasiparticles. The universality of fragmentation suggests potential for robust entanglement, though stability to environmental perturbations requires further study.
@@ Acknowledgements
The author thanks Olalla A. Castro-Alvaredo and M. A. Rajabpour for reading a previous versions of the draft and helpful comments and discussions. JZ acknowledgements support from the National Natural Science Foundation of China (NSFC) grant number 12205217 and Tianjin University Self-Innovation Fund Extreme Basic Research Project grant number 2025XJ21-0007.